# EFRame: Deeper Reasoning via Exploration-Filtering-Replay Reinforcement Learning Framework

### v1

@@ Abstract
Recent advances in reinforcement learning (RL) have significantly enhanced the reasoning capabilities of large language models (LLMs). Group Relative Policy Optimization (GRPO), an efficient variant of PPO that lowers RL’s computational cost, still faces limited exploration, low sample efficiency and instability, constraining its performance on complex reasoning tasks. To address these limitations, we introduce EFRame, an Exploration-Filtering-Replay framework that systematically augments GRPO along three critical dimensions. EFRame performs additional rollouts to explore high-quality trajectories, applies online filtering to eliminate low-quality samples that introduce noise and variance, and leverages experience replay to repeatedly exploit rare but informative samples. EFRame establishes a complete and stable learning cycle, guiding the model through a structured transition from exploration to convergence. Our experiments across a variety of reasoning benchmarks demonstrate that EFRame not only improves the robustness and efficiency of training, but also enables access to deeper reasoning capabilities that remain unattainable under vanilla GRPO. Furthermore, EFRame enables a more fine-grained categorization of training samples, allowing for a deeper analysis of how different types of samples contribute to the learning process in RL. Our code is avail able at https://github.com/597358816/EFRame.
@@ 1 Introduction
Reinforcement learning (RL) has played a pivotal role in the post-training of large language models (LLMs) GLM et al. (2024); Touvron et al. (2023). In the early stages, approaches such as Proximal Policy Optimization (PPO) and Direct Preference Optimization (DPO) were widely adopted under the Reinforcement Learning from Human Feedback (RLHF), effectively aligning model outputs with human preferences through preference-based rewards Schulman et al. (2017); Rafailov et al. (2023); Zhong et al. (2024); Wang et al. (2024b). More recently, the focus has shifted towards Reinforcement Learning with Verifiable Rewards (RLVR), which seeks to provide more consistent and scalable supervision by leveraging structured, verifiable reward signals Mroueh (2025). Within this paradigm, DeepSeek-R1 introduced Group Relative Policy Optimization (GRPO), a lightweight and efficient variant of PPO that compute relative rewards within rollout groups to eliminate the need for the critic model Shao et al. (2024); Liu et al. (2024); Guo et al. (2025). GRPO has demonstrated remarkable improvements in training efficiency and has shown promising potential in enhancing the reasoning capabilities of LLMs.
@@
RLVR relies on extracting reward signals from the model’s own pass@$k$ outputs to improve its pass@1 performance. However, this design inherently couples final performance to the choice of $k$, as the reward signal vanishes when all sampled responses are either correct or incorrect, making it impossible for the model to extract useful knowledge from such data. This constrains RLVR’s learning capacity to the base model’s pass@$k$, limiting further exploration and improvement Yue et al. (2025). Although increasing the sampling number $k$ per prompt may seem like a straightforward solution, it typically yields only marginal improvements while significantly reducing sample efficiency Lin et al. (2025). This results in GRPO exhibiting limited exploration and low sample efficiency, and indirectly leads to training instability in the later stages due to the disappearance of reward signals Yu et al. (2025); Li et al. (2025).
@@
To address the aforementioned issues, we propose EFRame: an Exploration-Filtering-Replay RL Framework. For challenging prompts that cannot be effectively handled by regular rollouts, EFRame introduces an additional rollout phase to enable targeted exploration. During this process, the model generates a small number of high-quality positive samples along with a large number of low-quality negative samples. To reduce training variance and improve efficiency, we apply an online filtering mechanism to discard low-quality samples. Meanwhile, high-advantage samples are stored in a replay buffer and reused during training to amplify their impact and enhance model stability.
@@
Previous studies have shown that both positive and negative samples contribute to RLVR training Zhu et al. (2025). With the support of the EFRame, we categorize training samples to better understand their roles in learning:
@@
  • High-quality samples consist of correct responses obtained through exploration on difficult prompts. They are crucial for activating deeper model knowledge and enhancing the performance ceiling, but their effective use requires experience replay to avoid instability caused by their low sampling probability.
@@
EFRame provides a systematic solution to the inherent trade-offs among training stability, efficiency, and exploration capacity in reinforcement learning. By enabling the model to effectively identify and leverage deeper latent knowledge, EFRame enhances the expressiveness and learning potential of large language models. Through the coordinated integration of targeted exploration, online sample filtering, and experience replay, the framework not only addresses the limitations of GRPO in handling complex reasoning tasks but also establishes a more principled and controllable training paradigm. Moreover, by introducing a categorization of training samples based on their difficulty and informativeness, EFRame offers a new analysis on the roles of different sample within the RLVR framework. Empirical evaluations confirm the efficacy of EFRame: on the challenging Geometry3K benchmark, it outperforms GRPO by 36.5%, and it achieves consistent gains across a wide range of multimodal and mathematical reasoning benchmarks, demonstrating strong generalizability and robustness.
@@ 2 Related Work
The early application of RL in LLMs and MLLMs OpenAI (2023); Team et al. (2024); Zhu et al. (2023); Wei et al. (2023); Liu et al. (2023) was primarily reflected in RLHF, where algorithms such as PPO were employed to align language model outputs with human preferences and Direct Preference Optimization (DPO) Rafailov et al. (2024) bypass the need for reward modeling and instead optimize the model policy directly from preference data. Recently, RL has advanced toward integrating rule-based reward mechanisms with sophisticated optimization techniques, as exemplified by models such as DeepSeek-R1 Guo et al. (2025) and Kimi-1.5 Team et al. (2025). As a lightweight variant of PPO, GRPO Shao et al. (2024) significantly improves training efficiency, but at the cost of increased instability. A series of subsequent studies have introduced optimizations from GRPO. These include gradient loss refinements, as demonstrated in DAPO Yu et al. (2025) and Dr.GRPO Liu et al. (2025); online filtering methods like Meng et al. (2025); advantage enhancement techniques, exemplified by VL-Rethinker Wang et al. (2025a) and Luffy Yan et al. (2025); clip acceleration methods such as CPPO Lin et al. (2025); curriculum learning strategies explored in Chen et al. (2025b; c); and unsupervised RL such as Wei et al. (2025a); Zhao et al. (2025); Agarwal et al. (2025). While existing methods typically focus on individual components, we systematically review the fundamental limitations of GRPO and propose a comprehensive framework with an integrated solution that addresses three critical aspects: exploration capacity, training stability, and sample efficiency. This helps create a more robust and effective GRPO-based training paradigm.
@@ 3.1 Gradient Loss
GRPO has been employed to enhance the reasoning capabilities of large language models, particularly in mathematical problem solving. Compared to PPO, GRPO improves computational efficiency by eliminating the need for a separate critic model. Instead, it estimates the advantage directly using group-normalized rewards. Given a question $q$ and a corresponding verifiable answer $a$, GRPO generates a set of responses $O=\{o_{i}\}_{i=1}^{G}$ using the policy model $\pi_{\theta_{old}}$, then optimizes the policy model by maximizing the following objective:
@@ 3.2 Limitations of GRPO
Instability. During the later stages of training, the GRPO algorithm may exhibit signs of instability or collapse, characterized by a sharp decline in reward and sudden spikes in both gradient and policy entropy, as illustrated in Fig. 2. In extreme cases, this can escalate into gradient explosion. Such instability presents a serious challenge to reliable long-horizon optimization, particularly in complex reasoning tasks where consistent gradient signals are critical for convergence. One contributing factor is the increased variance caused by the disappearance of relative advantages in the later stages of training. Supplementing with additional samples with non-zero advantages can help alleviate this issue.
@@ 4 Methodology
To address the issues in the previous section, we propose the Exploration-Filtering-Replay (EFR) framework, which consists of three main components: additional rollout, online filtering and replay buffer, as illustrated in Figure 3.
Figure 3: The overall workflow of EFR builds upon the vanilla GRPO framework by introducing three key components: additional rollout, online filtering, and replay buffer.
@@ 4.1 Regular Rollout
For each batch of prompts, we begin with a regular-rollout phase, following the original GRPO training process. This initial rollout serves two key purposes:
@@ 4.2 Additional Rollout
To address this, we introduce an additional rollout mechanism: we perform an extra sampling round (e.g. $G_{2}$=20, temperature = 1.2) for challanging prompts.
@@ 4.4 Replay buffer
While GRPO is fundamentally an online RL algorithm, we find that experience replay with high-quality offline samples remains highly effective. In particular, trajectories discovered during the second-round rollout—often of higher quality due to deeper exploration and carry rich reward signals. By selectively storing high-quality samples in the replay buffer, we amplify their influence during training, particularly when their gradients are frequently clipped due to low sampling probabilities. Furthermore, by enabling the model to repeatedly leverage these rare yet informative samples, experience replay plays a pivotal role in enhancing stability of entropy.
@@
This theorem was first introduced by Liu (2025), aand was organized and extended by cui2025entropymechanismreinforcementlearning. Proof can be seen in Liu (2025) and cui2025entropymechanismreinforcementlearning. ${H}$ indicates the policy entropy of policy model, and Cov denotes covariance, $\pi^{k}_{\theta}$ is the policy at step $k$, and $A^{\pi^{k}}(s,a)$ is the advantage function of action $a$ under state $s$. This result indicates that when high-advantage actions have low probability, they induce negative covariance, which increases entropy and reduce stability.
@@ 4.5 Discussion
The above subsections yield several practical insights regarding the interplay between Exploration-Filtering-Replay:
Additional rollout: By prompting LLM to generate multiple responses to challenging queries ($x\in\bar{\mathcal{D}}$) under a high temperature setting, the resulting positive samples are obtained with extremely low probability $\log\pi_{\theta}(y\mid x)$ and high advantage $\hat{A}(x,y)$. According Theorem 4.5, this high-advantage, low-probability scenario rapidly increases the policy entropy $H$, thereby preventing collapse of the entropy and maintaining effective exploration. At the same time, an excessive presence of low-quality samples can cause a nosedive of entropy, which is difficult to control. To address this, we filter out such samples and instead adopt a more controllable strategy to regulate entropy stability.
@@ 5.1 Experimental setup
Evaluation Benchmarks. Our experiments consist of two parts. First, we evaluate the overall performance of the algorithms on four open multimodal math benchmarks: MathVision Wang et al. (2024a), MathVista Lu et al. (2023), MathVerse Zhang et al. (2024), and We-Math Qiao et al. (2024). These benchmarks cover diverse problem types such as geometry, charts, and tables, spanning multiple subjects and levels of visual reasoning difficulty. Second, we conduct a detailed analysis for algorithms on the Geometry3K dataset to examine the algorithmic behaviors.
@@ 5.2 Main results
Performance. In our experiments, we employ the Exploration-Filtering-Replay framework, yielding notable improvements on both out-of-domain and in-domain datasets. As shown in Table 1, the proposed method delivers consistently strong results across four multimodal reasoning benchmarks, with an average gain of +5.66 over the base model (Qwen2.5-VL-7B-Instruct), outperforming GRPO by 19.9%. On the in-domain Geometry3K dataset, the improvement is even more pronounced: Table 2 reports a +16.80 gain, representing a 36.5% relative advantage over GRPO.
@@ 6.2 Online Filtering
Removing the online filtering module results in significant instability and inefficiency during training. Without filtering, the model treats all samples—regardless of their advantage—as equally important, allowing zero-advantage samples to increase gradient variance and low-quality negative trajectories to distorb the learning signal. This problem is further exacerbated during the additional rollout phase, where abundant incorrect samples significantly amplifies gradient variance, making it difficult for the model to retain meaningful patterns from rare high-quality responses. As a result, training becomes slower, more chaotic, and more prone to divergence.
@@
As shown in Fig. 5, during the early stages of training, the inclusion of a large number of negative samples significantly boosts the model’s performance on the training set, but yields little improvement on the test set. This indicates that most low-quality samples lead the model to converge more quickly to a local optimum, offering little benefit to its generalization ability. Moreover, the forward and backward computation of these negative samples incurs substantial overhead, reducing the sample efficiency of the overall training process. Online filtering plays a crucial role in maintaining gradient clarity and training stability by selectively preserving informative samples and discarding neutral or low-impact ones.
@@ 6.3 Replay Buffer
As high-quality samples typically have low generation probabilities, rewarding them shifts the model’s response distribution toward low-probability but correct outputs, disrupting the convergence trajectory and consequently increasing entropy. Due to their high advantage values and low sampling probabilities, high-quality samples are more likely to have their gradients clipped during training. In the absence of a replay buffer, such valuable samples contribute only short-lived and unstable gradients that are quickly overwritten. Instead of reinforcing beneficial behavior, this transient gradient signal increases policy uncertainty, pushing the model into a state of chaotic, high-entropy exploration, ultimately harming both learning stability and performance. To address this issue, we leverage experience replay to ensure that high-quality samples continuously provide effective gradients, enabling the model to converge along the new trajectory.

### v2

# EFRame: Deeper Reasoning via Exploration-Filter-Replay Reinforcement Learning Framework

@@ Abstract
Recent advances in reinforcement learning (RL) have significantly enhanced the reasoning capabilities of large language models (LLMs). Group Relative Policy Optimization (GRPO), an efficient variant of PPO that lowers RL’s computational cost, still faces limited exploration, low sample efficiency and instability, constraining its performance on complex reasoning tasks. To address these limitations, we introduce EFRame, an Exploration-Filter-Replay framework that systematically augments GRPO along three critical dimensions. EFRame performs additional rollouts to explore high-quality trajectories, applies online filtering to eliminate low-quality samples that introduce noise and variance, and leverages experience replay to repeatedly exploit rare but informative samples. EFRame establishes a complete and stable learning cycle, guiding the model through a structured transition from exploration to convergence. Our experiments across a variety of reasoning benchmarks demonstrate that EFRame not only improves the robustness and efficiency of training, but also enables access to deeper reasoning capabilities that remain unattainable under vanilla GRPO. Furthermore, EFRame enables a more fine-grained categorization of training samples, allowing for a deeper analysis of how different types of samples contribute to the learning process in RL. Our code is avail able at https://github.com/597358816/EFRame.
@@ 1 Introduction
Reinforcement learning (RL) has played a pivotal role in the post-training of large language models (LLMs) GLM et al. (2024); Touvron et al. (2023). In the early stages, approaches such as Proximal Policy Optimization (PPO) and Direct Preference Optimization (DPO) were widely adopted under the Reinforcement Learning from Human Feedback (RLHF), effectively aligning model outputs with human preferences through preference-based rewards Schulman et al. (2017); Rafailov et al. (2023); Zhong et al. (2024); Wang et al. (2024b). More recently, the focus has shifted towards Reinforcement Learning with Verifiable Rewards (RLVR), which seeks to provide more consistent and scalable supervision by leveraging structured, verifiable reward signals Mroueh (2025). Within this paradigm, DeepSeek-R1 introduced Group Relative Policy Optimization (GRPO), a lightweight and efficient variant of PPO that computes relative rewards within rollout groups to eliminate the need for the critic model Shao et al. (2024); Liu et al. (2024); Guo et al. (2025). GRPO has demonstrated remarkable improvements in training efficiency and has shown promising potential in enhancing the reasoning capabilities of LLMs.
@@
RLVR relies on extracting reward signals from the model’s own pass@$k$ outputs to improve its pass@1 performance. However, this design inherently couples final performance to the choice of $k$, as the reward signal vanishes when all sampled responses are either correct or incorrect, making it impossible for the model to extract useful knowledge from such data. This constrains RLVR’s learning capacity to the base model’s pass@$k$, limiting further exploration and improvement Yue et al. (2025). Although increasing the sampling number $k$ per prompt may seem like a straightforward solution, it typically yields only marginal improvements while significantly reducing sample efficiency Lin et al. (2025). This results in GRPO exhibiting limited exploration, and low sample efficiency, and indirectly leads to training instability in the later stages due to the disappearance of reward signals Yu et al. (2025); Li et al. (2025).
@@
To address the aforementioned issues, we propose EFRame: an Exploration-Filter-Replay RL Framework. For challenging prompts that cannot be effectively handled by regular rollouts, EFRame introduces an additional rollout phase to enable targeted exploration. During this process, the model generates a small number of high-quality positive samples along with a large number of low-quality negative samples. To reduce training variance and improve efficiency, we apply an online filtering mechanism to discard low-quality samples. Meanwhile, high-advantage samples are stored in a replay buffer and reused during training to amplify their impact and enhance model stability.
@@
Previous studies have shown that both positive and negative samples contribute to RLVR training Zhu et al. (2025). With the support of the EFRame, we categorize training samples to understand their roles in learning better:
@@
  • High-quality samples consist of correct responses obtained through the exploration of difficult prompts. They are crucial for activating deeper model knowledge and enhancing the performance ceiling, but their effective use requires experience replay to avoid instability caused by their low sampling probability.
@@
EFRame provides a systematic solution to the inherent trade-offs among training stability, efficiency, and exploration capacity in reinforcement learning. By enabling the model to effectively identify and leverage deeper latent knowledge, EFRame enhances the expressiveness and learning potential of large language models. Through the coordinated integration of targeted exploration, online sample filtering, and experience replay, the framework not only addresses the limitations of GRPO in handling complex reasoning tasks but also establishes a more principled and controllable training paradigm. Moreover, by introducing a categorization of training samples based on their difficulty and informativeness, EFRame offers a new analysis of the roles of different samples within the RLVR framework. Empirical evaluations confirm the efficacy of EFRame: on the challenging Geometry3K benchmark, it outperforms GRPO by 36.5%, and it achieves consistent gains across a wide range of multimodal and mathematical reasoning benchmarks, demonstrating strong generalizability and robustness.
@@ 2 Related Work
The early application of RL in LLMs and MLLMs OpenAI (2023); Team et al. (2024); Zhu et al. (2023); Wei et al. (2023); Liu et al. (2023) was primarily reflected in RLHF, where algorithms such as PPO were employed to align language model outputs with human preferences and Direct Preference Optimization (DPO) Rafailov et al. (2024) bypass the need for reward modeling and instead optimize the model policy directly from preference data. Recently, RL has advanced toward integrating rule-based reward mechanisms with sophisticated optimization techniques, as exemplified by models such as DeepSeek-R1 Guo et al. (2025) and Kimi-1.5 Team et al. (2025). As a lightweight variant of PPO, GRPO Shao et al. (2024) significantly improves training efficiency, but at the cost of increased instability. A series of subsequent studies have introduced optimizations from GRPO. These include gradient loss refinements, as demonstrated in DAPO Yu et al. (2025) and Dr.GRPO Liu et al. (2025); online filtering methods like Meng et al. (2025); advantage enhancement techniques, exemplified by VL-Rethinker Wang et al. (2025a) and Luffy Yan et al. (2025); clip acceleration methods such as CPPO Lin et al. (2025); curriculum learning strategies explored in Chen et al. (2025b; c); and unsupervised RL such as Zhang et al. (2025); Wei et al. (2025a); Zhao et al. (2025); Agarwal et al. (2025). While existing methods typically focus on individual components, we systematically review the fundamental limitations of GRPO and propose a comprehensive framework with an integrated solution that addresses three critical aspects: exploration capacity, training stability, and sample efficiency. This helps create a more robust and effective GRPO-based training paradigm.
@@ 3.1 Gradient Loss
GRPO has been employed to enhance the reasoning capabilities of large language models, particularly in mathematical problem-solving. Compared to PPO, GRPO improves computational efficiency by eliminating the need for a separate critic model. Instead, it estimates the advantage directly using group-normalized rewards. Given a question $q$ and a corresponding verifiable answer $a$, GRPO generates a set of responses $O=\{o_{i}\}_{i=1}^{G}$ using the policy model $\pi_{\theta_{old}}$, then optimizes the policy model by maximizing the following objective:
@@ 3.2 Limitations of GRPO
Instability. During the later stages of training, the GRPO algorithm may exhibit signs of instability or collapse, characterized by a sharp decline in reward and sudden spikes in both gradient and policy entropy, as illustrated in Fig. 2. In extreme cases, this can escalate into a gradient explosion. Such instability presents a serious challenge to reliable long-horizon optimization, particularly in complex reasoning tasks where consistent gradient signals are critical for convergence. One contributing factor is the increased variance caused by the disappearance of relative advantages in the later stages of training. Supplementing with additional samples with non-zero advantages can help alleviate this issue.
@@ 4 Methodology
To address the issues in the previous section, we propose the EFRame, which consists of three main components: additional rollout, online filtering and replay buffer, as illustrated in Figure 3.
Figure 3: The overall workflow of EFRame builds upon the vanilla GRPO framework by introducing three key components: additional rollout, online filtering, and replay buffer.
@@ 4.1 Regular Rollout
For each batch of prompts, we begin with a regular rollout phase, following the original GRPO training process. This initial rollout serves two key purposes:
@@ 4.2 Additional Rollout
To address this, we introduce an additional rollout mechanism: we perform an extra sampling round (e.g. $G_{2}$=20, temperature = 1.2) for challenging prompts.
@@ 4.4 Replay buffer
While GRPO is fundamentally an online RL algorithm, we find that experience replay with high-quality offline samples remains highly effective. In particular, trajectories discovered during the second-round rollout—are often of higher quality due to deeper exploration and carry rich reward signals. By selectively storing high-quality samples in the replay buffer, we amplify their influence during training, particularly when their gradients are frequently clipped due to low sampling probabilities. Furthermore, by enabling the model to repeatedly leverage these rare yet informative samples, experience replay plays a pivotal role in enhancing the stability of entropy.
@@
This theorem was first introduced by Liu (2025), and was organized and extended by Cui et al. (2025). Proof can be seen in Liu (2025) and Cui et al. (2025). ${H}$ indicates the policy entropy of policy model, and Cov denotes covariance, $\pi^{k}_{\theta}$ is the policy at step $k$, and $A^{\pi^{k}}(s,a)$ is the advantage function of action $a$ under state $s$. This result indicates that when high-advantage actions have low probability, they induce negative covariance, which increases entropy and reduces stability.
@@ 4.5 Discussion
The above subsections yield several practical insights regarding the interplay between Exploration-Filter-Replay:
Additional rollout: By prompting LLM to generate multiple responses to challenging queries ($x\in\bar{\mathcal{D}}$) under a high-temperature setting, the resulting positive samples are obtained with extremely low probability $\log\pi_{\theta}(y\mid x)$ and high advantage $\hat{A}(x,y)$. According to Theorem 4.5, this high-advantage, low-probability scenario rapidly increases the policy entropy $H$, thereby preventing the collapse of the entropy and maintaining effective exploration. At the same time, an excessive presence of low-quality samples can cause a nosedive of entropy, which is difficult to control. To address this, we filter out such samples and instead adopt a more controllable strategy to regulate entropy stability.
@@ 5.1 Experimental setup
Evaluation Benchmarks. Our experiments consist of two parts. First, we evaluate the overall performance of the algorithms on four open multimodal math benchmarks: MathVision Wang et al. (2024a), MathVista Lu et al. (2023), MathVerse Zhang et al. (2024), and We-Math Qiao et al. (2024). These benchmarks cover diverse problem types such as geometry, charts, and tables, spanning multiple subjects and levels of visual reasoning difficulty. Second, we conduct a detailed analysis of algorithms on the Geometry3K dataset to examine the algorithmic behaviors.
@@ 5.2 Main results
Performance. In our experiments, we employ the Exploration-Filter-Replay framework, yielding notable improvements on both out-of-domain and in-domain datasets. As shown in Table 1, the proposed method delivers consistently strong results across four multimodal reasoning benchmarks, with an average gain of +5.66 over the base model (Qwen2.5-VL-7B-Instruct), outperforming GRPO by 19.9%. On the in-domain Geometry3K dataset, the improvement is even more pronounced: Table 2 reports a +16.80 gain, representing a 36.5% relative advantage over GRPO.
@@ 6.2 Online Filtering
Removing the online filtering module results in significant instability and inefficiency during training. Without filtering, the model treats all samples—regardless of their advantage—as equally important, allowing zero-advantage samples to increase gradient variance and low-quality negative trajectories to distort the learning signal. This problem is further exacerbated during the additional rollout phase, where abundant incorrect samples significantly amplify gradient variance, making it difficult for the model to retain meaningful patterns from rare high-quality responses. As a result, training becomes slower, more chaotic, and more prone to divergence.
@@
As shown in Fig. 5, during the early stages of training, the inclusion of a large number of negative samples significantly boosts the model’s performance on the training set but yields little improvement on the test set. This indicates that most low-quality samples lead the model to converge more quickly to a local optimum, offering little benefit to its generalization ability. Moreover, the forward and backward computation of these negative samples incurs substantial overhead, reducing the sample efficiency of the overall training process. Online filtering plays a crucial role in maintaining gradient clarity and training stability by selectively preserving informative samples and discarding neutral or low-impact ones.
@@ 6.3 Replay Buffer
As high-quality samples typically have low generation probabilities, rewarding them shifts the model’s response distribution toward low probability but correct outputs, disrupting the convergence trajectory and consequently increasing entropy. Due to their high advantage values and low sampling probabilities, high-quality samples are more likely to have their gradients clipped during training. In the absence of a replay buffer, such valuable samples contribute only short-lived and unstable gradients that are quickly overwritten. Instead of reinforcing beneficial behavior, this transient gradient signal increases policy uncertainty, pushing the model into a state of chaotic, high-entropy exploration, ultimately harming both learning stability and performance. To address this issue, we leverage experience replay to ensure that high-quality samples continuously provide effective gradients, enabling the model to converge along the new trajectory.